# Optical computing using telecom device platform for nonlinear mapping to high-dimensional feature space

### v1

@@ Abstract
Mapping input signals to a high-dimensional space is a critical concept in various computing paradigms. We propose using commercially available telecom devices developed for high-speed optical data transmission for computing through nonlinear mapping of optical signals into a high-dimensional space where linear processing can be applied. We manipulate the output feature dimension by applying temporal up-sampling (at the speed of commercially available telecom devices) of input signals and a well-established wave-division-multiplexing (WDM). This gives remarkable flexibility in the dynamical phase masking of the input signal. We demonstrate the general concept employing readily available photonic devices, including a semiconductor optical amplifier and nonlinear optical loop mirror (equivalent nonlinear Mach-Zender interferometer). We apply linear regression on the high-dimensional output states to solve forecasting and classification problems. We investigate how nonlinear mapping provided by these devices can be characterized in terms of the increased controlled separability and predictability of the output state.
Keywords: Nonlinear mapping; Reservoir computing; Extreme learning machine; Nonlinear optical loop mirror; Optical computing
@@ 1 Introduction
Ultra-fast nonlinear photonic systems, in particular, are attractive for the implementation of unconventional computing approaches due to their relatively low power dissipation and capability of parallel signal processing. Recently, there has been a great deal of interest in the development of photonic-based ELMs and RCs (see, e.g., [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and references therein). Indeed, light possesses a rich set of degrees of freedom that can carry information. In optical communications, parameters such as amplitude, phase, wavelength, and polarization are routinely used to encode signals, with spatial division multiplexing emerging as an important technique for increasing data transmission rates further. Several examples of optical reservoir computing exploiting spectral signal multiplexing have already been explored [19, 20, 24, 25, 26, 27, 28]. In the temporal domain, up-sampling was used in [29]. However, the advantages of a faster output sampling rate were not leveraged because of the low bandwidth of the readout system. In this work, we demonstrate the advantage of up-sampling in enhancing the performance metrics of the ELM on different tasks. Also, we would like to point out that unlike the previously studied delay-based RC implementations, [29, 30], we propose approach without additional optical delay, while providing comparable processing capacity. Note that in the majority of the demonstrated optical reservoir computing and ELMs, only one of these degrees of freedom was used.
@@
In this work we propose and demonstrate how devices and technologies developed for optical data transmission can be used for computing applications. The combination of signal up-sampling in the temporal domain and WDM technology gives a great degree of flexibility in designing the structure of a high-dimensional output. Multiple spectral channels can range from coarse WDM ( systems with fewer than eight active wavelengths per fiber) to dense WDM (DWDM). DWDM can offer standard telecom solutions with a number of channels varying from tens to hundreds with typical (but also variable) channel spacing of 50GHz or 100GHz within the so-called optical fiber C-band (spectral interval from 1530 nm to 1565 nm). Traditional DWDM systems exploit wavelength-selective switches designed with fixed 50GHz or 100GHz filters. Using other fiber spectral bands, DWDM can be extended to thousands of channels. In the temporal domain, data streams with symbol rates as high as 32, 64 Gbaud (and more) in a single fiber can be produced with standard components. Thus, commercially available telecom devices can be utilized to produce a huge dimensional output feature space using only standard conventional technology. Numerous non-linear optical elements, modulators, devices and systems have been developed in the context of optical communications. To illustrate without loss of generality, the concept we consider here is a nonlinear optical loop mirror (NOLM) and semiconductor optical amplifier (SOA) as nonlinear transformers of optical signal.
@@ 2.2.1 Encoding techniques: dynamical masking
To feed information into the proposed optical ELM/RC for computing one needs to encode the input data stream in an electromagnetic field. Consider without loss of generality that the input data has a form of a vector $(s_{1},s_{2},...,s_{N})$. Encoding of the input messages (symbols) onto the optical waveform can be done in different ways exploiting the available degrees of freedom of light. The advantage of the proposed ELM/RC is that we can use easily accessible and well-developed telecom devices. For example, amplitude and/or phase modulation can be employed to encode input symbols in the amplitude/phase of the optical signal in the time domain. Encoding can be done in different frequency channels using established WDM technology. In addition to the symbol encoding, the effective dimensionality of the input can be increased via additional signal-invariant modulation, i.e., masking. Consider a WDM-based encoding scheme with $L$ spectral channels and an optical pulse train having the form:
$A(t)=\sqrt{P_{0}}\;\sum_{l=1}^{L}\sum_{j=1}^{J}x_{l,j}\;g(t-j\cdot T_{s})\exp\left(-i\omega_{l}t\right),$ (1)
@@
where $P_{0}$ is the power scaling parameter, $L$ is a number of spectral channels (indexed by $l$) used for coding of the information, $T_{s}$ is a symbol rate, $x_{l,j}$ for $j=1,2,\dots$ is a symbol (in general, a complex number) in the spectral channel $l$ at the temporal position $j$ that is used for encoding the input information in the optical domain. It can be sampled from a discrete set (alphabet) or be continuous. Function $g(t)$ describes the shape of a carrier pulse having the temporal scale $T_{0}$, which can be, in general, different from $T_{s}$. Thus, in the considered case, we need to encode the input data vector $(s_{1},s_{2},...,s_{N})$ into data array $x_{l,j}$. Evidently, this can be done in different ways, providing rich opportunities for the manipulation and optimization of the subsequent nonlinear mapping. Particular examples of encoding used here are described in section 2.5, while a comprehensive optimization of such coding will be presented elsewhere. To illustrate the general concept and without loss of generality, we consider here only simple intensity modulation ($x_{l,j}$ are real non-negative numbers). One can introduce asymmetry in the optical signal we use the skewed Gaussian pulse of the following form:
@@
Finally, the encoding mask can also be trained to optimize the performance of the ELM. In this case, we can tweak individual parameters of the encoding mask to maximize the accuracy of the proposed ELM. Figure 2 shows an example of an arbitrary encoding mask limited by the analog bandwidth of the arbitrary waveform generator.
@@ 2.3 Nonlinear transformation in optical domain
We consider here, without loss of generality, only two well-developed and easily accessible nonlinear devices that are widely used in telecom applications: (i) the nonlinear loop mirror (NOLM) [42], and (ii) semiconductor optical amplifier (SOA). NOLM operation can be explained as follows: the input signal power is divided between the two arms of a waveguide loop (for instance, an optical fiber), the signal phase in each arm is changed by the nonlinear propagation, and the resulting signal is formed by coupling the output ports of the arms (see next subsection for details). NOLM can produce overall nonlinear response using unequal coupling ratios, creating asymmetry of the accumulated nonlinear phase shifts, as in the original proposal [42] or by introducing imbalance in nonlinear propagation by using amplifiers (nonlinear amplifying loop mirror - NALM). Evidently, NOLM waveguide device can be realized on different material platforms. The same concept can be implemented in a nonlinear analog of the Mach–Zehnder interferometer, creating an interferometric phase converter to control the sign of the nonlinear phase shift [43].
SOA is a well-developed technology with many attractive characteristics, including compact size, efficient electrical pumping, cost-effectiveness and wideband gain [44, 45, 46]. However, in high-speed optical communication applications, the nonlinear properties of SOAs, relatively slow gain recovery time, and comparatively high noise figures (compared to other optical amplifiers) pose serious challenges. The carrier dynamics of SOAs have a characteristic scale of several hundred picoseconds. In the context of high-speed optical communications, this produces dependence of an instantaneous SOA gain on the input optical signal power that results in patterning effects - nonlinear distortions with memory. However, these nonlinear and inherent memory features can be attractive for optical computing applications, as demonstrated below. We describe the transfer functions for both NOLM and SOA below. However, without loss of generality, we present here results focusing on NOLM-based computing. Examples of optical computing using SOA-based high-dimensional mapping can be found in [47].
@@ 2.3.1 Nonlinear optical loop mirror
Figure 3: Input symbols, their optical encoding, NOLM output, and the samples captured by a photodetector.
@@ 2.3.2 Semiconductor optical amplifier
here in/out index denotes the input/output signal, $\beta$ is the linewidth enhancement (Henry) factor, $h_{0}$ parameter is related to the small signal gain $G_{0}=\exp(h_{0})$, $\tau_{c}$ is the gain recovery time, $E_{sat}$ is a characteristic saturation energy. It is seen that the nonlinear transformation of the input temporal signal stream by SOA creates an effective device memory defined by the delayed gain recovery: gain at a certain point in time $h(t)$ depends on the signal in the previous moments. We will demonstrate that this property can be exploited to create a high-dimensional feature output.
In the limit of short optical pulses, having temporal duration much smaller than $\tau_{c}$, we can use a well-known approximation [45] to simplify the description of SOA-induced transformation of the input signal:
[…]
This gives a simplified analytical description of the SOA-induced nonlinear transformation:
$P_{out}(t)$ $=P_{in}(t)\exp\left[h_{appr}(t)\right]$ (6)
$\phi_{out}(t)$ $=\phi_{in}(t)-\frac{\beta}{2}\,h_{appr}(t)$
@@ 2.4 Readout approaches to implement high-dimensional feature space
One of the key advantages of implementing nonlinear mapping to high-dimensional signals is the possibility to use simple processing, well-established telecom technologies and components at the readout layer. Though signal polarization and spatial modes can also be controlled and manipulated with telecom-grade devices, here, without loss of generality, we focus on the frequency and temporal domains at the output.
@@
Applying a modulator at the sampling rate, we can adjust the readout weight in the optical domain or, using high-speed optoelectronics available in telecom, change weights in the electrical domain.
In the frequency domain, the separation of different spectral channels can be done using well-developed WDM technology. The output optical signal passes through a WDM de-multiplexer that separates different spectral channels in the optical domain. The temporal up-sampling readout approach is applied to each spectral channel in parallel.
@@ 2.5 Training and validation
Here, $K$ denotes the dimensionality of the output vector, and the symbol $\rightarrow$ denotes the procedure of encoding the symbols in a waveform, propagating this waveform through the physical system and detecting it. The row vectors on the right-hand side of the arrows are the feature vectors that correspond to the input symbol sequences.
@@
In front of each (row) vector of features, we place the next symbol according to the sequence of symbols used to generate these features.
The process of training a model involves identifying the vector $\theta$ that minimizes the Mean Squared Error (MSE) between the predicted values, represented as $\mathbf{X}\theta$, and the actual target values, $\mathbf{Y}$. The goal is to find $\theta$ such that the difference between these predicted and actual values is as small as possible, which can be expressed as:
@@
$\min_{\theta}\|\mathbf{Y}-\mathbf{X}\theta\|^{2}_{2}+\lambda_{reg}\|\theta\|^{2}_{2}$ (12)
Here, $\lambda_{reg}$ is a regularization parameter that balances the trade-off between fitting the data well and keeping $\theta$ small. The adjusted singular values, denoted by $\mathbf{\Sigma}_{r}$, are defined as:
@@
We carefully adjusted $\lambda_{reg}$ to minimize NMSE on the validation subset.

### v2

# Optical neuromorphic computing via temporal up-sampling and trainable encoding on a telecom device platform

@@ Abstract
Mapping input signals to a high-dimensional space is a critical concept in various neuromorphic computing paradigms, including models such as Reservoir Computing (RC) and Extreme Learning Machines (ELM). We propose using commercially available telecom devices and technologies developed for high-speed optical data transmission to implement these models through nonlinear mapping of optical signals into a high-dimensional space where linear processing can be applied. We manipulate the output feature dimension by applying temporal up-sampling (at the speed of commercially available telecom devices) of input signals and a well-established wave-division-multiplexing (WDM). Our up-sampling approach utilizes a trainable encoding mask, where each input symbol is replaced with a structured sequence of masked symbols, effectively increasing the representational capacity of the feature space. This gives remarkable flexibility in the dynamical phase masking of the input signal. We demonstrate this approach in the context of RC and ELM, employing readily available photonic devices, including a semiconductor optical amplifier and nonlinear Mach-Zender interferometer (MZI). We investigate how nonlinear mapping provided by these devices can be characterized in terms of the increased controlled separability and predictability of the output state.
Keywords: Optical computing; Neuromorphic computing; Nonlinear mapping; Reservoir computing; Extreme learning machine
@@ 1 Introduction
Ultra-fast nonlinear photonic systems, in particular, are attractive for the implementation of unconventional computing approaches due to their relatively low power dissipation and capability of parallel signal processing. Recently, there has been a great deal of interest in the development of photonic-based ELMs and RCs (see, e.g., [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and references therein). Several notable works have demonstrated the potential of photonic reservoir computing in leveraging existing telecom technologies. A 16-node square mesh reservoir on a silicon photonics chip was implemented, capable of executing Boolean operations and header recognition tasks [24]. A parallel photonic reservoir computing approach using semiconductor optical amplifiers was also investigated, demonstrating competitive performance in speech recognition tasks [25]. A unified framework for reservoir computing and extreme learning machines was developed using a single nonlinear neuron with delayed feedback, realizable in optoelectronic and all-optical implementations [26]. Another study experimentally implemented reservoir computing with a nonlinear optoelectronic oscillator, achieving high performance on spoken digit recognition and time series prediction tasks [27]. A high-speed photonic reservoir computing system based on InGaAsP microring resonators was demonstrated for efficient all-optical pattern recognition in dispersive Fourier imaging [28].
Light possesses a rich set of degrees of freedom that can carry information. In optical communications, parameters such as amplitude, phase, wavelength, and polarization are routinely used to encode signals, with spatial division multiplexing emerging as an important technique for increasing data transmission rates further. Several examples of optical neuromorphic computing exploiting spectral signal multiplexing have already been explored, including implementations based on optical reservoir computing[19, 29, 30] as well as other optical neuromorphic architectures [20, 31, 32, 33]. Recent works demonstrated that WDM enhances photonic reservoir computing by enabling parallel processing and increasing computational capacity. A Fabry–Perot semiconductor laser-based RC leveraged multiple wavelength channels to improve signal equalization in optical communications [34]. A microring resonator-based RC exploited WDM for simultaneous multi-task processing, showing its potential for efficient parallel computing [35]. A waveguide-based RC demonstrated improved nonlinear signal equalization across multiple WDM channels [36]. Additionally, WDM has been shown to enhance the memory capacity of RC without requiring external optical feedback by using wavelength-multiplexed delayed inputs in microring-based architectures [37]. In this work, we show that WDM can also be used to enable a faster encoding mask, surpassing the speed limitations of single-channel implementations.
In the temporal domain, up-sampling was used in [38]. However, the advantages of a faster output sampling rate were not leveraged because of the low bandwidth of the readout system. In this work, we demonstrate the advantage of up-sampling in enhancing the performance metrics on different tasks. Also, we would like to point out that unlike the previously studied delay-based RC implementations, [38, 39], we propose approach without additional optical delay, while providing comparable processing capacity. Note that in the majority of the demonstrated optical reservoir computing and ELMs, only one of these degrees of freedom was used.
@@
In this work, we propose and demonstrate how devices and technologies developed for optical data transmission can be used for computing applications. While individual techniques like high-dimensional mapping, input masking, WDM, and specific nonlinear elements have been explored previously, our contribution focuses on a novel combination of these elements within a flexible framework designed for effective nonlinear mapping. The combined use of signal feature-space expansion via temporal up-sampling and WDM technology for parallel encoding gives a great degree of flexibility in designing the structure of a high-dimensional output. A key element of our approach is the incorporation of a trainable input encoding mask. This trainable mask provides a crucial advantage, allowing for task-specific optimization of the input signal representation before it undergoes nonlinear transformation, thereby enhancing the overall representational capacity and performance of the system. Multiple spectral channels can range from coarse WDM (systems with fewer than eight active wavelengths per fiber) to dense WDM (DWDM). DWDM can offer standard telecom solutions with a number of channels varying from tens to hundreds with typical (but also variable) channel spacing of 50GHz or 100GHz within the so-called optical fiber C-band (spectral interval from 1530 nm to 1565 nm). Traditional DWDM systems exploit wavelength-selective switches designed with fixed 50GHz or 100GHz filters. Using other fiber spectral bands, DWDM can be extended to thousands of channels. In the temporal domain, data streams with symbol rates as high as 32, 64 Gbaud (and more) in a single fiber can be produced with standard components. Thus, commercially available telecom devices can be utilized to produce a huge dimensional output feature space using only standard conventional technology. Numerous non-linear optical elements, modulators, devices and systems have been developed in the context of optical communications. To illustrate our combined approach featuring this trainable encoding, we consider here a balanced-arm MZI with non-symmetric couplers, which is mathematically equivalent to nonlinear optical loop mirror (NOLM), and semiconductor optical amplifier (SOA) as nonlinear transformers of optical signal. In what follows, we use balanced-arm MZI with non-symmetric couplers and NOLM interchangeably.
@@ 2.2.1 Encoding techniques: dynamical masking
To feed information into the proposed optical ELM/RC for computing one needs to encode the input data stream in an electromagnetic field. Consider that the input data has a form of a vector $(s_{1},s_{2},...,s_{N})$. Encoding of the input messages (symbols) onto the optical waveform can be done in different ways exploiting the available degrees of freedom of light. The advantage of the proposed ELM/RC is that we can use easily accessible and well-developed telecom devices. For example, amplitude and/or phase modulation can be employed to encode input symbols in the amplitude/phase of the optical signal in the time domain. Encoding can be done in different frequency channels using established WDM technology. In addition to the symbol encoding, the effective dimensionality of the input can be increased via additional signal-invariant modulation, i.e., masking. Consider a WDM-based encoding scheme with $L$ spectral channels and an optical pulse train having the form:
$A(t)=\sqrt{P_{0}}\;\sum_{l=1}^{L}\sum_{j=1}^{J}a_{l,j}\;g(t-j\cdot T_{s})\exp\left(-i\omega_{l}t\right),$ (1)
where $P_{0}$ is the power scaling parameter, $L$ is a number of spectral channels (indexed by $l$) used for coding of the information, $T_{s}$ is a symbol rate, $a_{l,j}$ for $j=1,2,\dots$ is a symbol (in general, a complex number) in the spectral channel $l$ at the temporal position $j$ that is used for encoding the input information in the optical domain. It can be sampled from a discrete set (alphabet) or be continuous.
@@
Function $g(t)$ describes the shape of a carrier pulse (or an encoding mask) having the temporal scale $T_{0}$, which can be, in general, different from $T_{s}$. Thus, in the considered case, we need to encode the input data vector $(s_{1},s_{2},...,s_{N})$ into array $a_{l,j}$. Evidently, this can be done in different ways, providing rich opportunities for the manipulation and optimization of the subsequent nonlinear mapping.For example, one can encode multiple consecutive symbols $s_{i}$ into multiple parallel instances of $a_{l,j}$, effectively parallelizing processing. Or one can introduce a time shift, when $a_{l,j}$ contain shifted copies of $s_{i}$ in different spectral channels, effectively introducing memory. To illustrate the general concept we consider here only simple intensity modulation ($a_{l,j}$ are real non-negative numbers).
In this work, spectral multiplexing is used differently depending on the task. For time-series forecasting, we introduce memory across channels by encoding time-delayed replicas of a base symbol sequence: specifically, we define $a_{l,j}=a_{1,j-l+1}$ for $l=1,\dots,L$ and $j\geq l$, such that channel $l$ contains the $(l-1)$-symbol delayed copy of the original sequence on channel 1. For classification tasks, we encode multiple features or input elements (image pixels) in parallel across different WDM channels at the same temporal index, effectively reducing the sequential processing length. These encoding schemes are shown in Tab. 1
Table 1: WDM-based encoding schemes. (a) Time-series forecasting uses delayed copies across spectral channels to introduce memory. (b) Classification encodes features in parallel across channels to reduce sequential depth.
(a) Time-series forecasting WDM Channel $j{=}1$ $j{=}2$ $j{=}3$ $j{=}4$ $j{=}5$.. $\omega_{1}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $\omega_{2}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $\omega_{3}$ $s_{1}$ $s_{2}$ $s_{3}$ $\omega_{4}$ $s_{1}$ $s_{2}$ $\omega_{5}$ $s_{1}$
(b) Classification WDM Channel $j{=}1$ $j{=}2$cdots $\omega_{1}$ $s_{1}$ $s_{6}$cdots $\omega_{2}$ $s_{2}$ $s_{7}$cdots $\omega_{3}$ $s_{3}$ $s_{8}$cdots $\omega_{4}$ $s_{4}$ $s_{9}$cdots $\omega_{5}$ $s_{5}$ $s_{10}$cdots
One can introduce asymmetry in the optical signal we use the skewed Gaussian pulse of the following form:
@@
Finally, the encoding mask can also be trained to optimize the performance of the ELM. In this case, we can tweak individual parameters of the encoding mask to maximize the accuracy of the proposed ELM. In this work, we used GWO [51] for global optimization of the encoding mask and Nelder-Mead simplex method for refinement. Figure 2 shows an example of an arbitrary encoding mask limited by the analog bandwidth of the arbitrary waveform generator.
@@ 2.3 Nonlinear transformation in optical domain
NOLM operation can be explained as follows: the input signal power is divided between the two arms of a waveguide loop (for instance, an optical fiber), the signal phase in each arm is changed by the nonlinear propagation, and the resulting signal is formed by coupling the output ports of the arms (see next subsection for details). NOLM can produce overall nonlinear response using unequal coupling ratios, creating asymmetry of the accumulated nonlinear phase shifts, as in the original proposal [52] or by introducing imbalance in nonlinear propagation by using amplifiers (nonlinear amplifying loop mirror - NALM). Evidently, NOLM waveguide device can be realized on different material platforms. The same concept can be implemented in a nonlinear analog of the Mach–Zehnder interferometer, creating an interferometric phase converter to control the sign of the nonlinear phase shift [53].
SOA is a well-developed technology with many attractive characteristics, including compact size, efficient electrical pumping, cost-effectiveness and wideband gain [54, 55, 56]. However, in high-speed optical communication applications, the nonlinear properties of SOAs, relatively slow gain recovery time, and comparatively high noise figures (compared to other optical amplifiers) pose serious challenges. The carrier dynamics of SOAs have a characteristic scale of several hundred picoseconds. In the context of high-speed optical communications, this produces dependence of an instantaneous SOA gain on the input optical signal power that results in patterning effects - nonlinear distortions with memory. However, these nonlinear and inherent memory features can be attractive for optical computing applications, as demonstrated below. We describe the transfer functions for both NOLM and SOA below. However, for clarity, we primarily present results based on NOLM-based computing. Examples of optical computing using SOA-based high-dimensional mapping can be found in [57].
The output state is obtained by taking the absolute value squared of the output of the complex transfer functions, i.e. we use intensity-only detection. We assume that the output signal is measured using photodetectors. Additionally, we consider the passive fiber losses are negligible compared to the signal modulation due to nonlinear signal transform.
@@ 2.3.1 Nonlinear optical loop mirror
Figure 3: Input symbols, their optical encoding, NOLM output, and the samples captured by a photodetector. For readability, 16 samples per symbol are shown instead of the 32 used in experiments.
@@
In this work, to calculate the input-output signal transformation in nonlinear MZI, we numerically simulate signal propagation trough fiber by solving Eq.(…) using the Fourth-Order Runge–Kutta in the Interaction Picture Method [58]. We do it for both arms, i.e. with initial conditions $\sqrt{\kappa}A_{\text{in}}(t)$ and $i\sqrt{1-\kappa}A_{\text{in}}(t)$. We then combine the solutions at $z=L_{\mathrm{NOLM}}$ as in Eq. (…) and take the absolute value squared to get the signal as registered by a photodetector.
@@ 2.3.2 Semiconductor optical amplifier
here in/out index denotes the input/output signal, $\beta$ is the linewidth enhancement (Henry) factor, $h_{0}$ parameter is related to the small signal gain $G_{0}=\exp(h_{0})$, $\tau_{c}$ is the gain recovery time, $E_{sat}$ is a characteristic saturation energy. It is seen that the nonlinear transformation of the input temporal signal stream by SOA creates an effective device memory defined by the delayed gain recovery: gain at a certain point in time $h(t)$ depends on the signal in the previous moments. In our previous work [57] we demonstrate that memory capacity associated with this effect to be $\geq 3.5$. We demonstrate that this property can be exploited to create a high-dimensional feature output. Eqs. 5 are directly integrated using Runge-Kutta method.
@@ 2.4 Readout approaches to implement high-dimensional feature space
One of the key advantages of implementing nonlinear mapping to high-dimensional signals is the possibility to use simple processing, well-established telecom technologies and components at the readout layer. Though signal polarization and spatial modes can also be controlled and manipulated with telecom-grade devices, here we focus on the frequency and temporal domains at the output.
@@
Applying a modulator at the sampling rate, we can adjust the readout weight in the optical domain or, using high-speed optoelectronics available in telecom, change weights in the electrical domain. Here, we simulate readout by multiplying the row-vector of intensities at the output of the ELM/RC with the regression column-vector, determined during the training procedure (see section 2.5).
In this work, we utilize WDM solely for encoding the input signal across multiple spectral channels. After the nonlinear transformation, these channels are not separated at the output stage. Instead, we assume a detection scheme where the combined optical output, containing all WDM channels, impinges on a single photodetector. The temporal up-sampling readout approach is then applied to this aggregated signal.
While this joint detection approach is employed here, it is worth noting that an alternative readout strategy could involve separating the different spectral channels using well-developed WDM technology. In such a scenario, the output optical signal would pass through a WDM de-multiplexer, and the temporal up-sampling readout approach could be applied independently to each spectral channel in parallel. However, exploring such parallel processing via output WDM separation is beyond the scope of the current study, which focuses on the impact of WDM encoding combined with joint detection.
@@ 2.5 Training and validation
Here, $K$ denotes the dimensionality of the output vector, $N$ denotes number of symbols fed into RC/ELM to generate features for next symbol prediction, $M$ denotes number of training samples and the symbol $\rightarrow$ denotes the procedure of encoding the symbols in a waveform, propagating this waveform through the physical system and detecting it. The row vectors on the right-hand side of the arrows are the feature vectors that correspond to the input symbol sequences.
@@
In front of each (row) vector of features, we place the next symbol according to the sequence of symbols used to generate these features. This is what we refer to as the single-step prediction task. We also evaluated our approach for autoregressive multi-step prediction, where the single-step prediction is applied iteratively, updating the time frame to incorporate newly predicted symbols.
The process of training a model involves identifying the regression vector $\theta$ that minimizes the Mean Squared Error (MSE) between the predicted values, represented as $\mathbf{X}\theta$, and the actual target values, $\mathbf{Y}$. The dimensionality of $\theta$ corresponds to the number of output features, $K$. In the following, we also use the Normalized Mean Squared Error (NMSE), defined as the MSE divided by the variance of $\mathbf{Y}$. NMSE provides interpretability, as $1-\text{NMSE}=R^{2}$, which represents the proportion of variance explained by our model. The goal is to find $\theta$ such that the difference between these predicted and actual values is as small as possible, which can be expressed as:
@@
$\min_{\theta}\|\mathbf{Y}-\mathbf{X}\theta\|^{2}_{2}+\lambda_{\text{reg}}\|\theta\|^{2}_{2}$ (11)
Here, $\lambda_{\text{reg}}$ is a regularization parameter that balances the trade-off between fitting the data well and keeping $\theta$ small. The adjusted singular values, denoted by $\mathbf{\Sigma}_{r}$, are defined as:
@@
We carefully adjusted $\lambda_{\text{reg}}$ to minimize MSE on the validation subset.
@@
2.6 Simulation hyperparameters
To ensure reproducibility and clarity, we summarize below all hyperparameters used across our numerical experiments in Table 2. These include system-specific parameters for the encoding and nonlinear transformation stages, as well as settings related to the learning process and readout. The table covers both nonlinear MZI/NOLM- and SOA-based implementations, along with details of the receiver and training configuration.
Category Parameter Value
Encoding Symbol interval $T_{s}$ 400 ps
Samples per Symbol 32
Peak power $P_{0}$ 50 W
Num WDM channels 5
WDMrequency step $\Delta\nu$ 50 GHz
Encoding mask size 8
Analog bandwidth 20 GHz
MZI/NOLM Nonlinearity $\gamma$ $0.8\;\text{W}^{-1}\text{km}^{-1}$
GVD $\beta_{2}$ $26$ $\text{ps}^{2}\text{km}^{-1}$
Length L 100 m
Coupling ratio $\kappa$ 0.3
SOA Recovery time $\tau$ 200 ps
Henry factor $\beta$ 5
Log gain $h_{0}$ 6.91
Saturation energy $E_{\text{sat}}$ 8 pJ
Receiver Sampling rate 80 GSa/s
Samples per symbol 32
Learning Training symbols 5000
Testing symbols 100-5000
Symbols $N$ 15
Regularization $\lambda_{\text{reg}}$ (optimized per task and device SOA/NOLM) $10^{-8}$ to $10^{-2}$
Table 2: All hyperparameters used in the simulation.